OpenStudy (anonymous):

4/2 = d^3 solve for d - please show working

OpenStudy (anonymous):

For this you have to ask your self what "undoes" the cube function

OpenStudy (anonymous):

ok, i understand if you take a square over it becomes a square root on other side right. x^2 = 1 x=sqrt1 right?

OpenStudy (anonymous):

is it d = 3sqrt(4/2)

OpenStudy (anonymous):

Correct, you got it!

OpenStudy (anonymous):

thanks mate, its late and im doing a structural engineering subject, just wanted to check, all the best.

OpenStudy (anonymous):

one note: because a negative number times a negative number equals a positive, namely, \[(-a)^2=a^2\], you would have \[x^2=4 \implies x=+ or - 2\] but this is not the case with odd numbered powers, i bet you could work this out too see why

OpenStudy (anonymous):

real and complex numbers is that right?

OpenStudy (anonymous):

no its not that complicated, it just has to do with what happens when you multiply a negaitive by a negative. \[-3*-3*-3 = -27, while 3*3*3 = 27, but -3*-3 = 9= 3*3\]